OpenStudy (anonymous):

have a hard time with math all together. I question is w=C * 3,963square -2

OpenStudy (anonymous):

Is the equation: a) \(w = C * 3963^2 -2\) b) \(w = C * 3963^{-2}\) c) \(w = (C * 3963)^2 -2\) d) \(w = (C * 3963)^{-2}\) e) None of the above?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Ok, so First thing to simplify is to distribute that exponent. \[(a^1*b^1)^c = a^{1*c} * b^{1*c} \]

OpenStudy (anonymous):

Polpak popcorn is back whenever you get time. :-)

OpenStudy (anonymous):

\(C = C^1 \) and \(3963 = 3963^1\) right?

OpenStudy (anonymous):

Shouldn't have used c in my example. But the point being you multiply the exponents.

OpenStudy (anonymous):

So what does your expression simplify to when you distribute the -2 exponent?

OpenStudy (anonymous):

does the negatuve exponent change or stay the same

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

my problem now is D=1.2\[\sqrt{h}\]